OpenStudy (anonymous):

IQ scores (as measured by the Stanford-Binet intelligence test) are normally distributed with a mean of 100 and a standard deviation of 16. Find the approximate number of people in the U.S. (assuming a total population of 280,000,000) with an IQ higher than 110. Round your answer to the nearest 100,000.

OpenStudy (anonymous):

@douglaswinslowcooper @Hero @SithsAndGiggles @whpalmer4 PLEASE HELP

OpenStudy (anonymous):

z = (value - mean)/ (standard deviation) use cumulative normal distribution to find fraction with IQ larger than corresponding to this z-score. If z = 1, 16% have higher values, so (0.16)(280 x 10^6) or 44, 800,000. For your question value= 110, mean = 100, SD = 16 {I thought it was 15, check that] z = (110-100/16 = etc. then look it up in tables or on net

OpenStudy (anonymous):

it would then be 1-.63 right?

OpenStudy (anonymous):

or no not .63 but .7357

OpenStudy (anonymous):

If your z-score indicates the fraction to be 0.63, then 1-.63 would be above, yes. Multiply that by the total population.. But your z-score is 0.63, which the tables for normal distribution indicate 0.736 is the fraction with z<= 0.63. Use 0.736 times total population, as I see you now realize.

OpenStudy (anonymous):

i got 206,080,000 which is wrong by the key

OpenStudy (anonymous):

perhaps "nearest 100,000" is key, thus 206,100,000

OpenStudy (anonymous):

in the key it says the answer is 74,500,000.. is there perhaps something else I'm leaving out?

OpenStudy (anonymous):

I would have done z = (110-100)/15 = 0.667 p = 0.745 1-p = .255 (0.255) (280x10^6) = 71,400,000 I think this is right.

OpenStudy (anonymous):

so do you think he maybe made a typo? because in the key the answer is 74,500,000

OpenStudy (anonymous):

all i want to know is how to do it right because i have an exam tomorrow